Problem #2

Refrigerant 134a enters an adiabatic compressor at 0.1 MPa and -20 oC while it leaves at 1 MPa and 110 oC. The mass flow rate of the compressor is 0.3 kg/s. Determine: a) The power input to the compressor, in kW. b) The volume flow rate of the refrigerant at the compressor inlet.
Answer #1

h inlet = 395.6 kJ/kg

h outlet = 493.7 kJ/kg

Enthalpy values obtained from p-h curve of R134a

Mass flow =0.3 kg/sec

(A)

Power input = m*(hout - hin) = .3*(493.7 - 395.6) =29.43KWatt

(B)

Specific volume at compressor inlet

= 0.2074 (m^{3})/kg (obtained form p-h curve of R134-a)

Any doubt are welcome in commentbox

Please upvote and rate positive

Thanks in advance

Please upvote